OpenStudy (banjoman):

The sum of the ages of Corey and Karissa is 75 years. 12 years ago, Corey's age was 2 times Karissa's age. How old is Corey now? will medal for answer!!

jhonyy9 (jhonyy9):

C + K = 75 2K +K = 75+12 =>3K =87 => K = 23 so than you know that K=23 now just substitute this value of K inside this equation of C+K=75 and will get right easy the Corey"s age solve these equations hope helped